### v1

# YourMT3+: Multi-instrument Music Transcription with Enhanced Transformer Architectures and Cross-dataset Stem Augmentation

Abstract
Multi-instrument music transcription aims to convert polyphonic music recordings into musical scores assigned to each instrument. This task is challenging for modeling as it requires simultaneously identifying multiple instruments and transcribing their pitch and precise timing, and the lack of fully annotated data adds to the training difficulties. This paper introduces YourMT3+, a suite of models for enhanced multi-instrument music transcription based on the recent language token decoding approach of MT3. We strengthen its encoder by adopting a hierarchical attention transformer in the time-frequency domain and integrating a mixture of experts (MoE). To address data limitations, we introduce a new multi-channel decoding method for training with incomplete annotations and propose intra- and cross-stem augmentation for dataset mixing. Our experiments demonstrate direct vocal transcription capabilities, eliminating the need for voice separation pre-processors. Benchmarks across ten public datasets show our models’ competitiveness with, or superiority to, existing transcription models. Further testing on pop music recordings highlights the limitations of current models. Fully reproducible code and datasets are available at https://github.com/mimbres/YourMT3.
Index Terms—  Multi-instrument, automatic music transcription (AMT), music information retrieval (MIR), transformers, data augmentation, mixture of experts (MoE), music tokens
1 Introduction
Automatic music transcription (AMT) [1] is a fundamental task in music information retrieval where the goal is to transform music audio input into a sequence of musical notes, with each note possessing properties such as onset, offset, pitch, and sometimes velocity. The output is typically presented in the form of MIDI or piano-roll notation. The significance of AMT extends to a wide range of applications, including interactive music systems [2], automatic accompaniment generation [3], and music performance assessment.
The key challenge of this research is multi-instrument AMT: identification and transcription of various instruments with vocals from music recordings. Recently, there has been notable progress in this field: MT3 [4] utilized a MIDI-like decoding transformer, while PerceiverTF [5] employed a spectral attention transformer that generates conventional piano-roll. Unfortunately, the absence of fully reproducible code for these models has been a significant limitation for replication and further research. Our replication of MT3, trainable from scratch, is dubbed as YourMT3 [6]. Based on this, we propose YourMT3+, a hybrid architecture that incorporates advanced architectures and training methods for further enhancements. YourMT3+ and its variants differ from prior works [4, 5] in the following key aspects:
  • Encoder: PerceiverTF [5], which generated piano-rolls, is now trained with the MT3 framework to generate note event tokens. We replaced MT3’s encoder with PerceiverTF featuring spectral cross attention (SCA). Additionally, replacing its feedforward network (FFN) with a mixture of experts (MoE) [7], denoted as YPTF.MoE, demonstrates promising results.
  • Decoder: Tokens for singing transcription have been further defined. We introduce a multi-channel decoder that replaces MT3’s single-channel decoder [4]. This enables task-query based training and the use of partially annotated data, improving performance.
  • Augmentation: The proposed online data augmentation framework incorporates intra-stem and cross-stem mixing across datasets and pitch-shifting. In particular, cross-stem augmentation allows for transcribing singing with other instruments without the need for a voice separation front-end.
  • Evaluation: Our models were extensively validated on various multi-instrument and single-instrument datasets. One of the main applications of multi-instrument AMT can be transcribing pop music. We provide refined annotations for the existing pop music dataset [8], presenting the first study to investigate multi-instrument AMT performance on commercial pop music.
2 Relation to Prior Work
While substantial research exists in AMT, multi-instrument transcription has recently seen significant developments. The field often faces challenges due to the scarcity of fully annotated datasets for all instruments, making it low-resourced. Strategies such as multi-task learning [4, 9], unsupervised learning methods [10] and iterative re-alignment techniques [11] have offered partial remedies, with most models producing piano-roll outputs at the frame level.
Compared to the conventional AMT models based on onsets and frames [12], MT3 [4] is a sequence-to-sequence model that mainly distinguished itself in decoding outputs. It decodes a note-level representation similar to language tokens derived from MIDI, deviating from the traditional frame-level piano-rolls. In Section 3.3, we discuss the advantages of using these output tokens in YourMT3.
The transcription of singing within multi-instrument AMT remains largely unexplored, despite potential overlaps with source separation [13] and melody extraction [14]. PerceiverTF [5], a model with piano-roll output, has significantly advanced the transcription of multiple instruments and vocals by introducing spectral cross-attention (SCA) and stem dataset mixing. We propose an augmentation method, denoted by a plus (+) sign, that formalizes the earlier stem mixing approach [5] within an online multi-dataset pipeline.
Fig. 1: Overview of YourMT3+. (left) Our encoder $E(\cdot)$ takes as input a log mel spectrogram $S$ derived from audio $X$. (center) An auto-regressive decoder $D(\cdot)$ with the language model (LM) head is conditioned by $E(S)$, and output event tokens $Y^{\prime}$. (right) Cross-dataset stem augmentation, described in Section 4.
3 Model
In the YourMT3+ model taxonomy, we designate models with the same architecture and training method as MT3 as YMT3. Models that replace the encoder with PerceiverTF, trained using MT3’s single-channel decoder and our stem augmentation, are termed YPTF+Single. The variant replacing the encoder’s FFN with a mixture of experts (MOE) is called YPTF.MoE. These models have demonstrated that the PerceiverTF encoder can efficiently process language-style tokens, such as MIDI, instead of using piano-roll. Additionally, we propose a multi-channel decoder that assigns instrument groups per channel and masks loss for unannotated instruments, allowing training with incomplete labels. The final model, which integrates all these features, is termed YPTF.MoE+Multi.
The left panel of Figure 1 provides a detailed overview of our final extended model, YPTF.MoE+Multi. The subsequent subsections will detail the components of our model variants, including the audio input, encoder, decoder, and output tokens.
3.1 Input
[…]
3.2 Encoder
The encoder $E(\cdot)$ takes $S$ as an input, where the last dimension of $S$ typically matches the encoder’s hidden dimension $d$. Our baseline encoder of YMT3 is based on the T5-small v1.1 [15] encoder composed of 8 standard transformer blocks with 6-head self-attention and gated FFNs. The proposed YPTF replaces the encoder with PerceiverTF (PTF) [5] blocks as depicted in Figure 1 (left).
[…]
MoE: Our models, designated as YPTF.MoE, replace the FFNs in latent and temporal transformer blocks with mixture of experts (MoE) layers from [7], routing attention output to two out of eight experts (see Supplemental B.5). In our experiments, MoE increased the model complexity by about 5% while improving performance across various datasets. Unlike PerceiverTF, we use RoPE [17] in every sub-block of the encoder to integrate positional information through rotation matrices, replacing trainable position embedding (PE), and pre-LayerNorm with pre-RMSNorm. However, these modifications only offered minor benefits in memory and computation without significantly impacting performance.
3.3 Output Tokens
The center panel of Figure 1 shows the output sequence $Y^{\prime}$ with a maximum $N$ time steps, and the tokens representing MIDI-like events are listed in Section F of Supplemental Document. We followed the note sequence structure proposed in MT3 [4], and two modifications were made to the MT3 tokens: (a) unused velocity tokens except 0 and 1 were removed, and (b) programs 100 and 101 were reserved for singing voice (melody) and singing voice (chorus), respectively.
Compared to traditional piano-roll [12, 18, 10, 9, 5], the use of MIDI-like tokens [4] offers several advantages. First, it is more memory-efficient, using tokens for note onset, shift, and offset rather than hundreds of frames for silence as in piano-rolls. This efficiency also simplifies handling of multi-instrument data; piano-rolls require separate matrices for each instrument group, while the token approach needs only a program token for each change. Second, using tokens allows expansion of program vocabulary without a substantial memory growth, unlike piano-rolls. Lastly, tokens explicitly represent note onsets and offsets, avoiding the extra post-processing required for piano-rolls.
3.4 Decoder
We use an auto-regressive decoder $D(\cdot)$, conditioned on the encoder’s last hidden state, to generate note sequences. The baseline decoder, based on T5-small v1.1 and denoted as Single, produces a single sequence with events from multiple instruments.
When annotations are available for only one or some instruments in the audio, we need to mask the loss for unannotated instruments. The Single decoder’s output blends multiple programs, making it hard to mask specific instruments due to token dependencies. To address this, we propose a Multi decoder. It can provide separately maskable supervision for each latent $L$ of the PTF encoder, allocated into channels for each program group.
In our implementation, the PTF encoder’s output hidden states are grouped by allocating two latents per channel. With group-linear projection, $k=26$ latents result in $k^{\prime}=13$ projected channels. The Multi decoder then independently decodes each of the $k^{\prime}$ inputs, producing $k^{\prime}$ sequences for each program using parallel decoders with shared parameters. We set the maximum sequence length to $N_{\text{single}}=1024$ (as in MT3 [4]) and $N_{\text{multi}}=256$. Potential truncation loss is discussed further in Supplemental B.6.
4 Data Augmentation
This section describes an augmentation method for training with multiple datasets. Our strategy is to maximize the diversity of the training examples by randomly mixing selected stems from across multiple datasets. Intra-stem augmentation described in Section 4.1 involves selectively muting stems within a multi-track recording to generate several variations, as demonstrated with MT3 [4] and the Slakh dataset. The concept of cross-dataset stem augmentation, as discussed in Section 4.2, draws inspiration from PerceiverTF [5]. It aims to create a new mixture of stems from multiple datasets. Additionally, we employ pitch-shifting as described in Section 4.3.
4.1 Intra-stem Augmentation
This refers to the process of randomly dropping instruments from a segment containing multiple stems. From any dataset we sample $X$, a 2.048-second segment starting from a random point. Assuming that $X$ is composed of $N$ stems denoted $x_{1},x_{2},\ldots,x_{N}$, we define a set $\hat{X}_{\text{in}}$ of randomly selected or dropped stems as:
$\hat{X}_{\text{in}}=\{x_{i}:x_{i}\in X,\text{ with }x_{i}\sim\text{Bernoulli}(p)\}$ (1)
with $i\in\{1,2,...,N\}$ where $N>1$. Here, $p$ represents the probability of each stem being selected, set to 0.7 as default. Each $x_{i}$ may or may not be chosen with a probability $p$, resulting in $\hat{X}_{\text{in}}$ containing various combinations of stems from the original segment $X$.
Algorithm 1 Cross-dataset Stem Augmentation
0:  $X$, $U$, $L$, $J$, $\Psi$, $\tau$, $p$ $X$: A segment $X\in U$, with stems $x\in X$. $U$: Cached segment batches from various datasets. $L$: Maximum length of sequence. 1,024 by default. $J$: Maximum number of iterations w.r.t $j$. 5 by default. $\Psi$: Stem mixing policy. $\tau$: Exponential decay parameter. 0.3 by default. $p$: Probability for intra stem selection. 0.7 by default. }
1:  $\hat{X}_{\text{in}}\leftarrow{x_{i}:x_{i}\in X,\text{selected with }x_{i}\sim\text{Bernoulli}(p)}$
2:  $\hat{X}_{\text{ex}}\leftarrow\emptyset$
3:  $j\leftarrow 0$
4:  while $r\sim\textit{Uniform}(0,1)<e^{-\tau j}\text{ and }|\hat{X}_{\text{ex}}|<L$ and $j<J$ do
5:     $X^{\prime}\leftarrow$ a randomly sampled segment from $U\setminus X$
6:     $X^{\prime}\leftarrow\text{Filter}(X^{\prime};\Psi)$ // retain stems meeting criteria
7:
8:     if $X^{\prime}\neq\emptyset$ then
9:        $\hat{X}_{\text{ex}}\leftarrow\hat{X}_{\text{ex}}\cup{X^{\prime}}$ // add stems
10:        $j\leftarrow j+1$
11:     end if
12:  end while
13:  $\hat{X}\leftarrow\hat{X}_{\text{in}}\cup\hat{X}_{\text{ex}}$
14:  $\text{Mix}(\hat{X})$ // apply stem mixing
4.2 Cross-dataset Stem Augmentation
Procedure: In Algorithm 1, we designate $U$ as a collection of cached segment batches across diverse datasets, with its size required to be at least equal to the batch size and preferably larger, if permitted by memory constraints. The base segment $X$ is a sampled segment from $U$, and the elements of $X$ are stems denoted by $x$. Here, $x$ signifies a stem ID, including related token and audio information.
Intra-stem augmentation is first applied to $X$ as in Equation 1, yielding a processed base segment $\hat{X}_{\text{in}}$. Next, we enter a loop to mix the base stems of $\hat{X}_{\text{in}}$ with the stems coming from other segments. $U\setminus X$ represents the set of all segments in $U$ excluding $X$. Each iteration begins by randomly sampling a segment $X^{\prime}$ from $U\setminus X$. Stems in $X^{\prime}$ that do not satisfy policy $\Psi$ (detailed in Supplemental D.2) are then filtered out. Subsequently, $\hat{X}_{\text{ex}}$ is updated by merging $X^{\prime}$. This loop persists until at least one stopping criterion described in the following subsection is satisfied. Once the aggregation is complete, the Mix($\cdot$) function executes the actual mixing of tokens and audio content in a batch-wise manner.
Stopping criteria In Line 5 of Algorithm 1, three criteria are established to stop the iterative mixing among stems. The first criterion is an exponential decay $S(j)$ that serves as the survival function defined as $S(j)=e^{-\tau j},$ where $\tau$ controls the surviving curve with respect to $j$-th iteration. The second criterion restricts $\hat{X}_{\text{ex}}$ to a length $L$, measured as sequence length post-tokenisation. The last criterion, $j>J$ with $J=5$ allows mixing up to 5 segments per base segment.
4.3 Pitch-shifting
During training, pitch-shifting is applied to a batch, after applying cross-dataset stem augmentation. The phase vocoder-based pitch-shifting on the GPU was adapted from the TorchAudio [^1] https://pytorch.org/audio library. We used the default settings with the exception of setting $\textit{nFFT}=512$ for time-stretching. All the elements within the batch are randomly assigned to one of five groups, each with an equal chance of selection. Each group is then subjected to a pitch shift by an interval of -2, -1, 0, +1, or +2 semitones.
5 Experiments
Train Test
MusicNet-EM, GuitarSet, MIR-ST500, ENST-Drums, Slakh, EGMD, Maestro, CMedia, URMP, SMT-Bass MusicNet, MusicNet-EM, GuitarSet,MIR-ST500, ENST-Drums, Slakh, Maestro, MAPS, URMP, RWC-Pop (refined)
Table 1: Summary of datasets for train/test. Multi-instrument datasets with full annotation and stems are highlighted in light blue, while those with partially annotated instruments are highlighted in pink. (refined) We offer updated annotations for RWC-Pop [8].
5.1 Experimental Setup
Data Preparation: Table 1 lists the datasets used for training and evaluating our model. We offer a software package, [an anonymised link], for dataset setup and split information to ensure reproducibility of our results. Audio data was converted into 16 kHz mono WAV format. Stems were stored as arrays, and mix-tracks as WAV files, also treating stemless tracks as mix-tracks. For training our Single decoder models on MIR-ST500 [19] and CMedia [20] , we produced singing and accompaniment stems using a pre-trained separation model [13]. With the Multi decoder, we also incorporated the original mix tracks from these datasets.
Evaluation Metrics: To evaluate transcription accuracy for each instrument, we employ the Instrument Note Onset F1 metric  [5]. This metric, valid for any instruments including drums, requires matching the onset, pitch, and instrument to the reference within a tolerance of ±50 ms. For multiple non-drum instruments, we additionally utilize the Instrument-Agnostic Onset F1 and Offset F1 necessitating exact matches for only onset or both onset and offset. These metrics parallel the standard Note F1 metrics [21] for single-instrument datasets. Furthermore, we used the Multi (instrument offset) F1 metric [4] for evaluating multi-instrument AMT systems, where correct predictions require matching onset-offset pairs, pitch, and instrument type, excluding drum offsets. Our Multi F1 metric is notably more stringent than the Multi Onset F1 reported for PerceiverTF [5].
Vocabulary: Our models were trained using MT3_FULL_PLUS and tested on MT3_MIDI_PLUS, detailed in Section F of the Supplemental Document. Despite testing exclusively with the MIDI vocabulary, results in Table 3, labeled +full vocab, show that training with the more fine-grained FULL vocabulary enhanced performance compared to training and testing solely with MIDI.
Test Set Instrument YMT3 YMT3+ YPTF+S YPTF+M YPTF.MoE+M MT3 [4] AMT
noPS noPS |||| PS noPS |||| PS noPS |||| PS noPS |||| PS (colab) Baseline *
MAPS [22] (unseen) Piano 81.44 85.92 |||| 87.73 88.37 |||| 88.73 87.84 |||| 86.88 87.88 |||| 86.25 80.62 88.40 [23$\clubsuit$
MAPS [22] (seen) - - - - - - 85.14 [24$\clubsuit$
Maestro v3 94.78 94.80 |||| 94.31 96.28 |||| 95.85 95.59 |||| 94.54 96.98 |||| 96.52 94.86 97.44 [24$\clubsuit$
MusicNet ext.
(EM) [11]
Strings 81.69 89.04 |||| 88.34 88.39 |||| 89.39 88.52 |||| 87.04 91.32 |||| 90.07 $\triangle$ 80.00 [11] $\sharp$
Winds 74.95 82.91 |||| 80.53 77.72 |||| 79.59 77.18 |||| 76.54 83.46 |||| 78.50 $\triangle$ 85.50 [11] $\sharp$
MusicNet ext.
 [10, 11]
Strings 58.20 64.67 |||| 63.94 64.63 |||| 65.40 64.17 |||| 64.08 66.14 |||| 66.09 $\triangle$ 63.90 [11] $\sharp$
Winds 50.76 55.58 |||| 55.05 52.55 |||| 54.27 51.82 |||| 51.42 55.95 |||| 55.33 $\triangle$ 60.90 [11] $\sharp$
MIR-ST500 [19] (SVS) Singing 67.98 70.39 |||| 70.69 70.82 |||| 70.56 71.07 |||| 71.32 71.60 |||| 72.05 $\lozenge$ 70.73 [25]
MIR-ST500 [19] 3.62 64.03 |||| 65.69 66.75 |||| 67.11 69.67 |||| 70.26 70.59 |||| 71.07 $\lozenge$ 78.50 [5]
MIR-ST500 (100ms [20]) 3.64 71.15 |||| 72.08 73.26 |||| 73.89 79.29|||| 80.63 81.14|||| 82.08 $\lozenge$ -
ENSTdrums (DTP [26]) Drums 87.77 87.60 |||| 87.40 89.72 |||| 90.65 88.68 |||| 90.61 88.79 |||| 89.48 77.82 84.50 [26] $\clubsuit$
ENSTdrums (DTM [26]) 78.64 81.84 |||| 83.09 85.65 |||| 86.41 85.14 |||| 87.18 85.92 |||| 87.27 70.31 79.00 [26] $\clubsuit$
GuitarSet [27] (MT3 [4]) Guitar 88.53 91.39 |||| 88.49 91.61 |||| 88.32 88.92 |||| 86.74 91.65 |||| 88.87 89.10 91.10 [5]
URMP [28] Onset F1 [4] Agnostic 77.10 80.00 |||| 81.47 81.11 |||| 81.54 74.56 |||| 75.72 81.05 |||| 81.79 76.65 77.0 [4]
URMP [28] Multi F1 [4] Ensemble 58.23 62.13 |||| 62.03 64.34 |||| 65.89 57.25 |||| 59.82 67.22 |||| 67.98 58.71 59.0 [4]
Slakh [29] Onset F1 [4] Agnostic 64.83 77.96 |||| 75.28 80.70 |||| 76.32 79.39 |||| 75.68 84.14 |||| 84.56 75.20 81.9 [5]
Slakh [29] Multi F1 [4] All 61.77 65.92 |||| 63.61 69.52 |||| 65.13 69.37 |||| 64.96 73.98 |||| 74.84 57.69 62.0 [4$\heartsuit$
Table 2: Dataset-wise Note Onset F1. PS and noPS represent training with and without pitch shifting augmentation, respectively. (EM) denotes evaluation using refined labels [11]. (SVS) refers to experiments using singing separated audio as input, obtained through Spleeter [13]. (DTP) represents using drum and percussion as input. (DTM) uses input including drum, percussion, and accompaniment. The Onset F1 score on Slakh is instrument-agnostic F1 for non-drum classes. ($\triangle$) Unavailable due to training split overlaps. ($\clubsuit$) Single-instrument AMT. ($\lozenge$) Singing voice class was not defined. ($\sharp$) Additionally collected synthetic data from 8.5K songs were used for pre-training [11].
Training: Our models were trained with two NVIDIA A100 GPUs using BFloat16 mixed-precision. In the implemented online data pipeline, four CPU processes per GPU were allocated to efficiently load and augment data without causing streaming bottlenecks. In our preliminary experiments, we tested three optimizers at a constant learning rate of 1e-03: AdaFactor [30], AdamW [31], and AdamWScale [32]. AdamWScale, a variant of AdamW that normalizes gradients using root-mean-square (RMS) energy, provided the most efficient training. Our models were trained using AdamWScale and a cosine scheduler for 300K steps, with initial and final learning rates of [1e-02, 1e-05] and a 1,000-step warm-up from 1e-03. We set the dropout rate at 0.05.
Model Onset F1 Offset F1 Drum F1
YMT3 base 64.8 41.7 77.8
   + Intra-aug. +4.8 +5.5 +0.6
   + Full-vocab. +0.6 +2.1 +2.6
   + Data balancing + 4.0 +4.7 +1.3
   + Cross-aug. +4.0 +7.2 +1.6
   + PTF-encoder +1.8 +4.2 +1.9
   + FFN $\rightarrow$ MoE +1.5 +1.3 +3.7
   + Multi decoder +1.8 +4.0 +0.6
YPTF.MoE+Multi 84.6 70.7 90.1
MT3 (colab) 75.2 56.8 83.9
MT3 [4] 76 57 -
PerceiverTF [5] 81.9 - 78.3
Table 3: Model component analysis and comparison on the Slakh [29] dataset. (-) Values not reported.
Fig. 2: Instrument Onset F1 on Slakh [29].
Fig. 3: Instrument-Onset/Frame/Multi F1 on RWC-Pop [8].
5.2 Results and Discussion
In Table 2, our models are compared with other state-of-the-art models across datasets. From MAPS to GuitarSet, evaluations use Instrument Note Onset F1, while URMP and Slakh are assessed using Instrument-agnostic Note Onset F1 and Multi F1. Due to space constraints, only the top-performing baselines (*) are listed on the table’s rightmost column. Details of all models are available in our project repository.
Our models prefixed by Y- outperformed MT3 [4] across all datasets. Notably, our models and the unseen baseline [23], trained without MAPS [22], outperformed the baseline [24] trained on MAPS. This is likely due to the Maestro [33] dataset being about nine times larger, providing significantly more in-domain knowledge. Among our models, YPTF.MoE+Multi matched or exceeded the performance of the latest baseline models in most datasets. It showed exceptional performance on both refined and unrefined datasets in MusicNet strings, particularly in tests with refined labels (EM [11]). However, a noticeable under-performance was observed in singing transcription compared to the baseline [5]. As evidenced by about 10% higher F1 on the MIR-ST500 (100ms), many onset timing errors exceeded the acceptable 50ms range and fell within 100ms. Given that our model and the baseline [5] share similar encoder structures, our decoder may be more prone to timing errors than traditional piano-roll models. Additionally, the practicality of a 100ms onset tolerance, used in past MIREX [20] singing transcription protocol, appears justified.
YMT3+ and YPTF+Single differ only in their encoders. This comparison revealed that the PTF encoder architecture performs particularly well in complex multi-instrument datasets such as MIR-ST500, ENSTdrums (DTM), and Slakh. Cross-stem augmentation, denoted by the (+) symbol in model names, proved essential for transcribing singing without singing voice separation (SVS). YMT3 recorded an F1 score of 3.6% without separation, while YMT3+ with augmentation reached 64%. The models with Multi decoders were beneficial when training on partially annotated datasets, such as MIR-ST500 and ENSTdrums. Mixture of Experts (MoE) showed consistent performance improvements across all datasets. Notably, while pitch-shifting often led to performance degradation in other models, YPTF.MoE compensated for this loss or even improved performance, as evidenced by the Slakh result.
As compared in the lower section of Table 2, YPTF.MoE+ Multi significantly outperformed the baselines (MT3 [4] and PerceiverTF [5]) on multi-instrument datasets such as URMP and Slakh. The baseline Multi F1 score marked with a $\heartsuit$ is from MT3 authors’ report [4]. For the complete comparison table with MT3 [4] and PerceiverTF [5], see Section H of the Supplemental Document.
Ablation Study: In Table 3, the impact of each model component on performance was investigated. Both intra-and cross-stem augmentations significantly improved performance by over 4 percentage points, while all other proposed components steadily enhanced transcription performance. Additionally, the performance improvement denoted by Data balancing suggested that previously adopted temperature-based sampling in MT3 [4] might not be suitable for determining the sampling probability of AMT datasets. This is further discussed in Section F of the Supplemental Document.
Performance on Pop Music: As seen at the bottom of Table 3, our model demonstrated competitive performance on the synthetic dataset [29] compared to other multi-AMT models. In Figure 3, our final model achieved 50 to 90% performance for most instruments, except for a few non-mainstream ones like chromatic percussion (c. perc) and synth pad (s.pad) in the synthetic dataset. However, it significantly underperformed on commercial pop music recordings as shown in Figure 3. Particularly for non-main instruments (excluding piano, bass, vocals, and drums), our models performed below 10%. This indicates that relying solely on synthetic datasets is insufficient for modeling the diverse timbres of pop music. Furthermore, except for the piano, all the pitched instruments showed a significant gap in the chroma-level metric, suggesting substantial octave errors and hinting that more varied pitch-shifting could be beneficial.
6 Conclusion and Future Work
This work presented YourMT3+, a hybrid model suite that combines MT3 and PerceiverTF features. Our final model, YPTF.MoE +Multi, employed spectral cross-attention and a Mixture of Experts in its encoder for enhanced performance, and a multi-channel decoder to handle the instruments where annotation is partially available. Our models trained using the proposed online augmentation strategy demonstrated direct vocal transcription capabilities without the need for a singing separation front-end. The final model significantly outperformed MT3 and PerceiverTF on the multi-AMT benchmark with a parameter increase of less than 2.5% compared to MT3. Evaluations across ten public datasets also validated our model’s competitiveness. Despite progress, challenges persist: onset timing in singing voice transcription lags behind our baseline, and low performance in pop music may stem from reliance on synthetic datasets for diverse instruments. Future research will address these issues.